OpenStudy (anonymous):

Evaluate: sin10+sin20+...+sin90

OpenStudy (anonymous):

is it not witht the squares

OpenStudy (anonymous):

i mean we shud use cos^2 x=1-sin^2x

OpenStudy (anonymous):

nt with square....it ws a mistake

OpenStudy (tkhunny):

Are those supposed to be degrees? If you really mean 10, rather than 10º, that is quite a different matter.

OpenStudy (john_es):

Let θ be then angle expressed in radians. \[\sum_{i=1}^{9}\sin(n\theta)=\frac{\sin\theta-\sin(10\theta)+\sin(9\theta)}{2-2\cos\theta}=6.21503\]